# K-theory and localizing invariants of large categories

### v2

@@ 4.2. Localizing invariants of dualizable categories
The notion of a localizing invariant is defined in the same way as for By Theorem C.6 the category is -presentable, and again we will consider only accessible localizing invariants. It turns out that any localizing invariant defined on small categories has a unique extension to a localizing invariant defined on dualizable categories, so that
Roughly the idea is the following. Proposition 1.62 shows that for a compactly generated presentable category the Calkin category is, up to direct summands, the quotient of by the full subcategory of compact objects For a general dualizable category one can think of as a “virtual quotient of by compact objects”. We will use this point of view to extend localizing invariants from to
@@
By Propositions 4.6 and 4.8, the assignment is indeed a functor By Proposition 4.7, it is a right inverse to the precomposition
On the other hand, for any and any dualizable category we have a short exact sequence
@@
Further, we have two straightforward results about commutation of with sufficiently nice pullbacks and pushouts.
@@
We observe that there is an alternative equivalent definition of as follows.
@@ 4.3. Right Kan extension
Let be an accessible localizing invariant, where is a presentable stable category. Then the functor is the right Kan extension of via
We will again use the fact that the category is -presentable (Theorem C.6).
It is convenient to define the following topologies on the opposite categories and In both cases we define a pretopology for which a family of functors is a cover if consists of one element and the functor is fully faithful. To deal with set-theoretic issues we also consider the induced pretopologies on the small subcategories and (note that these subcategories are closed under finite limits), where is an uncountable regular cardinal. In each case we denote by the corresponding Grothendieck topology.
We make the following observation.
@@
It suffices to prove this for Let be a fully faithful strongly continuous functor between dualizable categories. Consider the Čech conerve of i.e. – the cofiber product of copies of It follows from Proposition 4.12 that is the Čech conerve of Since is stable, we conclude that ∎
Next, we observe the following equivalence of categories of sheaves.
Proposition 4.18.
Let be an uncountable regular cardinal, and let be a presentable category. Let be a -sheaf. Then the composition is also a -sheaf. The induced functor
is an equivalence. Its inverse is given by the right Kan extension.
Proof.
The first statement is obvious. The functor sends covers to covers, and it commutes with finite limits. It remains to apply Theorem C.6, part 3), condition (iv): it implies that for any fully faithful strongly continuous functor between -compact dualizable categories we can find a fully faithful strongly continuous functor where is -compact. In particular, the covering of is a refinement of the covering This implies the equivalence of categories of sheaves, as stated.∎
@@
Choose an uncountable regular cardinal such that and commute with -filtered colimits. It suffices to show that the restriction of to is the right Kan extension of the restriction of to This follows immediately from Propositions 4.17 and 4.18: and are -sheaves respectively on and and is the composition ∎
@@ 4.4. Example: sheaves on the real line
We consider the special case of sheaves on the real line. Let be a dualizable category. Recall that the category is dualizable [Lur18]. In this subsection we give a detailed proof of the following special case of Theorem 0.2.
@@
Recall that if is compactly generated, then we have a notion of a singular support for -valued sheaves on -manifolds [RS18, Definition 4.5]. We will be interested in the following special case.
Denote by the cone of non-positive integers. The -topology on is defined as follows: a subset is -open if is open and In other words, is -open if either or or for some We denote by the set equipped with -topology. Denote by the identity map of considered as a map of topological spaces. Similarly, we denote by the map to equipped with -topology. The following is due to Kashiwara and Schapira.
@@
In fact there is a reasonable notion of singular support for sheaves with values in arbitrary presentable stable categories, see Remarks 4.24 and 4.25. To avoid the technicalities, we will use the following definition in the special case
Notation. For a presentable stable category we denote by resp. the category resp. of -valued sheaves on resp.
Note that for we have a tautological equivalence where the latter category was defined in Subsection 1.6. Here the poset is considered as a compactly assembled category. In particular, by Proposition 1.35 the category is dualizable.
Moreover, by Remark 1.36 for any presentable stable category we have
@@
Let be a dualizable category, and choose some dense subset Denote by the set considered as a linearly ordered set with the usual order.
1) We have a short exact sequence in
Here the functor is given by
2) We have
@@
Thus, and the maps are zero for In particular, is strongly continuous and the adjunction counit is an equivalence, so is a localization.
Now, we observe that the category is naturally equivalent to the right orthogonal This equivalence sends a sheaf to the functor Note that for we have hence
Using (4.6), we see that the category consists of functors such that for any the map is an isomorphism. This corresponds exactly to the fact that is a sheaf on
Finally, since the functor has a right adjoint, we have equivalences
This proves that we have a short exact sequence (4.4) in This proves part 1).
2) By (4.3), it is sufficient to prove the vanishing for Take and consider the short exact sequence (4.4). We need to show that the map is an equivalence in
Note that the categories and are compactly generated. We have an -indexed semi-orthogonal decomposition where the (fully faithful, strongly continuous) functors are defined in (4.5). Also, we have a tautological -indexed orthogonal decomposition of the category The functor is compatible with these decompositions and it induces the identity functors on the components. Hence, the map is an isomorphism. ∎
@@
where the is the inclusion. Hence, it suffices to prove the result for sheaves on
We claim that there is a natural short exact sequence in
Assuming this, the result follows from Proposition 4.22.
The functor in (4.7) is the left adjoint to The functor is given by where is the map (extending ) to the sobrification of Note that as a set is naturally identified with (where is the generic point), and clearly
An explicit description of the functors and is the following. For a convex open subset and for we have
@@
If is a -manifold and is a presentable compactly generated stable category, then the singular support of a -valued sheaf on is defined in [RS18, Definition 4.5], which is a straightforward generalization of [KS90, Definition 4.4.1]. However, if is a dualizable category which is not compactly generated, then already in the -dimensional case the non-characteristic deformation lemma [RS18, Theorem 4.1] fails and the conditions of [KS90, Proposition 5.1.1] are no longer equivalent.
To illustrate this, let Consider a -valued sheaf on given by for Then according to [RS18, Definition 4.5] the singular support would be the zero section But clearly the sheaf is not locally constant, hence the definition of the singular support has to be modified.
We mention several equivalent ways to define the singular support for a sheaf on a -manifold with values in a presentable stable category
1) A point is not contained in if and only if for some (any) choice of local coordinates it satisfies the equivalent conditions and of [KS90, Proposition 5.1.1].
2) The singular support is the smallest conical closed subset such that is contained in the essential image of the functor
This functor is fully faithful since the inclusion has a left adjoint.
3) The singular support of is the support of the conical sheaf (of spectra) on (microlocal Hom), where the bifunctor is defined as in [KS90, Definition 4.4.1].
4) Choose a regular cardinal such that takes values in Then the composition defines a sheaf with values in the compactly generated category (since the Yoneda embedding commutes with limits). We put where the latter singular support is defined as in [RS18, Definition 4.5].
5) If is dualizable, we can define the singular support of to be the singular support of the sheafification of the presheaf where the latter is defined as in [RS18, Definition 4.5].
@@ 4.5. Commutation of K-theory with infinite products
We first recall a simple proof of the commutation of with infinite products.
Proposition 4.27 (Heller’s criterion).
[KW19, Lemma 2.12] Let be a small triangulated category, and take two objects The following are equivalent:
@@
1) Let be an accessible stable category with a non-degenerate -structure. Let be accessible localizing invariants and a morphism such that is an isomorphism. Then is an isomorphism.
2) In particular, both -theory and continuous -theory commute with infinite products.
@@
By Proposition 1.103 it suffices to prove the statement for continuous -theory. Consider a collection of dualizable categories. We need to show that We may assume that the categories are all equivalent: indeed, each is a retract in of the category
In other words, we need to show that for a set and a dualizable category the map is an isomorphism. But this is the special case of 1) for the standard -structure on Indeed, by Proposition 1.105 the functor is a localizing invariant, and so is the functor When is compactly generated, the map is an isomorphism by Corollary 4.28.∎
@@
Put The functors are compatible with the functors hence we obtain a functor By construction, we have We need to show that in
For each the inclusions define a fully faithful functor Its left adjoint is simply the structural functor Denote by the composition functors, Since the compositions vanish for we see that the direct sum of the functors is well-defined. Moreover, we have
Now, denote by the composition, and put (where ). Equivalently, is the semi-orthogonal projection onto the left orthogonal to (where ). Denote by the natural transformations, Then
Since the composition vanishes for we see that the direct sum is also a well-defined endofunctor of By the above, we obtain an exact triangle
@@
Since we may assume that Applying 4.22 to we obtain a short exact sequence
in Moreover, the source and the target of are compactly generated categories. Denote by the induced functor on the subcategories of compact objects. It suffices to show that is a -equivalence. As in the proof of proposition 4.22, we put and
Note that the functor has an obvious section which corresponds to the collection of “representable” functors where is defined in (4.5). We need to show that in
Choose any bijection Denote by the full stable subcategory generated by the objects Then is equivalent to the category where as usual denotes the poset Clearly, each inclusion has both left and right adjoints. Moreover, we see that the elements form a basis of
It follows from Corollary 4.31 that Since we deduce that
@@
is also an isomorphism. By induction, we see that is an isomorphism for
Denote by the induced map between the localizing invariants of dualizable categories. Arguing as above, we see that is an isomorphism for It remains to show that if is an isomorphism for some then is also an isomorphism.
Let be a dualizable category. Then the functor is a strongly continuous localization (the right adjoint sends an object of to the corresponding constant sheaf). We denote by the kernel of Applying Proposition 4.32, we see that the composition
@@ 5.1. Continuous posets
Recall that a partially ordered set is called continuous if has directed supremums (joins), and the functor has a left adjoint Equivalently, this means the following.
Assuming that has directed supremums, one defines another relation on namely, if for any directed system of elements of such that we have for some Then is continuous if and only if for each the poset is directed and we have equality
@@ 5.2. Sheaves of spectra on continuous posets
Let be a continuous poset. We denote by the category of sheaves on (where is considered as a category) for the above topology. This is in fact the category from Subsection 1.6.
Namely, since every object of has the smallest covering sieve, a presheaf is a sheaf if and only if
@@ 5.3. Localizing invariants of categories of sheaves on continuous posets
Consider now a presheaf on with values in for We denote by the restriction functors for and by their right adjoints. Consider the category of -valued presheaves on (i.e. the category of sections of the associated biCartesian fibration over ). For we denote by the functor given by
The right adjoint to is given by
Denote by the category of -valued sheaves on For we have a functor The right adjoint to is given by
Our goal in this section is to prove the following result.
@@
1) The category is dualizable.
2) For each compact element the functor is strongly continuous and fully faithful.
3) The functors induce an isomorphism
@@
In particular, we have for This shows that for the functor commutes with coproducts. This proves 2).
To prove 1), we first observe that the category is dualizable. Indeed, we have a -indexed semi-othogonal decomposition of the category in
Since are dualizable, so is the category by Proposition 1.82.
Now, it follows from (5.2) that the sheafification functor has a left adjoint We deduce that the category is dualizable.
To prove 3), consider the full subcategory By the above we have a short exact sequence in
where the functor is the left adjoint to the inclusion. We first observe the following.
Claim 1. Let be a presheaf such that for any the map is zero. is Then is contained in
@@
Next, we observe that the category has the following semi-orthogonal decomposition.
Claim 2. For the functor is zero (resp. fully faithful) for (resp. for ). Moreover, we have a semi-orthogonal decomposition
@@
Indeed, it follows from (5.2) (by adjunction) that for the object is left orthogonal to On the other hand, it follows from Claim 1 that the object is contained in (since the latter presheaf can take non-zero values only for such that but ). This proves the isomorphism (5.5).
We deduce that
@@
Hence, the adjunction unit is an isomorphism. We conclude that the functor is fully faithful for
Similarly, if and then for This implies the semi-orthogonality.
Finally, taking into account the vanishing of for we deduce that the categories generate This proves the claim.∎

### v3

@@ 4.2. Localizing invariants of dualizable categories
The notion of a localizing invariant is defined in the same way as for By Ramzi’s theorem [Ram24a] (see also Theorem C.6 below), the category is -presentable, and again we will consider only accessible localizing invariants. It turns out that any localizing invariant defined on small categories has a unique extension to a localizing invariant defined on dualizable categories, so that Roughly the idea is the following. Proposition 1.61 shows that for a compactly generated presentable category the Calkin category is, up to direct summands, the quotient of by the full subcategory of compact objects For a general dualizable category one can think of as a “virtual quotient of by compact objects”. We will use this point of view to extend localizing invariants from to
@@
By Propositions 4.6 and 4.8, the assignment is indeed a functor By Proposition 4.7, it is a right inverse to the precomposition On the other hand, for any and any dualizable category we have a short exact sequence
@@
Further, we have two straightforward results about commutation of localizing invariants with sufficiently nice pullbacks and pushouts in
@@
We observe that there is an alternative equivalent way to extend localizing invariants to dualizable categories.
@@ 4.3. Right Kan extension
Let be an accessible localizing invariant, where is an accessible stable category. Then the functor is the right Kan extension of via
We will again use the fact that the category is -presentable (Ramzi’s theorem), and also one of the characterizations of -compact objects in (for ), given by Theorem C.6. It is convenient to define the following topologies on the opposite categories and In both cases we define a pretopology for which a family of functors is a cover of if consists of one element and the functor is fully faithful. To deal with set-theoretic issues we also consider the induced pretopologies on the small subcategories and (note that these subcategories are closed under finite limits), where is an uncountable regular cardinal. In each case we denote by the corresponding Grothendieck topology. Note that in each case is a subcanonical topology. Although the category is not necessarily complete (it is only assumed to be accessible and stable), we still have a notion of an -valued sheaf. Namely, the sheaf condition only requires the existence of the relevant limits, not all limits. We make the following observation.
@@
It suffices to prove this for Let be a fully faithful strongly continuous functor between dualizable categories. Consider the Čech conerve of i.e. – the cofiber product of copies of It follows from Proposition 4.12 that is the Čech conerve of Since is stable, we conclude that More precisely, for any morphism in its Čech conerve has the following property: for the map is an isomorphism. ∎
@@
Choose an uncountable regular cardinal such that is -accessible, and and commute with -filtered colimits. It suffices to show that the restriction of to is the right Kan extension of the restriction of to Let be a -compact dualizable category. By Theorem C.6, we can find a fully faithful strongly continuous functor such that is compactly generated and is -compact in Denote by the Čech conerve of in Then each is a compactly generated category, and is -compact. Now, the functor is a sheaf of spaces, covered by the representable sheaf with the Čech nerve By Proposition 4.17, the functors and are sheaves with values in It follows that the right Kan extension of from to exists, and its value on is given by
@@ 4.4. Example: sheaves on the real line
As an example, we will compute the finitary localizing invariants of the category of sheaves on the real line. Let be a dualizable category. Recall that the category is dualizable [Lur18]. We will give a detailed proof of the following special case of Theorem 0.2.
@@
Recall that if is compactly generated, then we have a notion of a singular support for -valued sheaves on -manifolds [RS18, Definition 4.5]. We will be interested in the following special case. Denote by the cone of non-positive real numbers. The -topology on is defined as follows: a subset is -open if is open and In other words, is -open if either or or for some We denote by the set equipped with -topology. Denote by the identity map of considered as a map of topological spaces. Similarly, we denote by the map to equipped with -topology. The following is due to Kashiwara and Schapira.
@@
In fact there is a reasonable notion of singular support for sheaves with values in arbitrary presentable stable categories, see Remarks 4.23 and 4.24. To avoid the technicalities, we will use the following definition in the special case Notation. For a presentable stable category we denote by resp. the category resp. of -valued sheaves on resp. Note that for we have a tautological equivalence where the latter category was defined in Subsection 1.6. Here the poset is considered as a compactly assembled category. In particular, by Proposition 1.34 the category is dualizable. Moreover, by Remark 1.35 for any presentable stable category we have
@@
Let be a dualizable category, and choose some dense subset Denote by the set considered as a linearly ordered set with the usual order. 1) We have a short exact sequence in
Here the functor is given by 2) We have
@@
Thus, and the maps are zero for In particular, is strongly continuous and the adjunction counit is an equivalence, so is a localization. Now, we observe that the category is naturally equivalent to the right orthogonal This equivalence sends a sheaf to the functor Note that for we have hence
Using (4.6), we see that the category consists of functors such that for any the map is an isomorphism. This corresponds exactly to the fact that is a sheaf on Finally, since the functor has a right adjoint, we have equivalences
This proves that we have a short exact sequence (4.4) in This proves part 1). 2) By (4.3), it is sufficient to prove the vanishing for Take and consider the short exact sequence (4.4). We need to show that the map is an equivalence in Note that the categories and are compactly generated. We have an -indexed semi-orthogonal decomposition where the (fully faithful, strongly continuous) functors are defined in (4.5). Also, we have a tautological -indexed orthogonal decomposition of the category The functor is compatible with these decompositions and it induces the identity functors on the components. Hence, the map is an isomorphism. ∎
@@
where the is the inclusion. Hence, it suffices to prove the result for sheaves on We claim that there is a natural short exact sequence in
Assuming this, the result follows from Proposition 4.21. The functor in (4.7) is the left adjoint to The functor is given by where is the map (extending ) to the sobrification of Note that as a set is naturally identified with (where is the generic point), and clearly An explicit description of the functors and is the following. For a convex open subset and for we have
@@
If is a -manifold and is a compactly generated presentable stable category, then the singular support of a -valued sheaf on is defined in [RS18, Definition 4.5], which is a straightforward generalization of [KS90, Definition 4.4.1]. However, if is a dualizable category which is not compactly generated, then already in the -dimensional case the non-characteristic deformation lemma [RS18, Theorem 4.1] fails and the conditions of [KS90, Proposition 5.1.1] are no longer equivalent. To illustrate this, let Consider a -valued sheaf on given by for Then according to [RS18, Definition 4.5] the singular support would be the zero section But clearly the sheaf is not locally constant, hence the definition of the singular support has to be modified. We mention several equivalent ways to define the singular support for a sheaf on a -manifold with values in a presentable stable category 1) A point is not contained in if and only if for some (any) choice of local coordinates it satisfies the equivalent conditions and of [KS90, Proposition 5.1.1]. 2) The singular support is the smallest conical closed subset such that is contained in the essential image of the functor
This functor is fully faithful since the inclusion has a left adjoint. 3) The singular support of is the support of the conical sheaf (of spectra) on (microlocal Hom), where the bifunctor is defined as in [KS90, Definition 4.4.1]. 4) Choose a regular cardinal such that takes values in Then the composition defines a sheaf with values in the compactly generated category (since the Yoneda embedding commutes with existing limits). We put where the latter singular support is defined as in [RS18, Definition 4.5]. 5) If is dualizable, we can define the singular support of to be the singular support of the sheafification of the presheaf where the latter is defined as in [RS18, Definition 4.5].
@@ 4.5. Commutation of K-theory with infinite products
We first recall the argument showing the commutation of of small triangulated categories with infinite products.
Proposition 4.26.
[KW19, Lemma 2.12](Heller’s criterion) Let be a small triangulated category, and take two objects The following are equivalent:
@@
1) Let be an accessible stable category with a non-degenerate -structure. Let be accessible localizing invariants and a morphism such that is an isomorphism. Then is an isomorphism. 2) In particular, both -theory and continuous -theory commute with infinite products.
@@
By Proposition 1.103 it suffices to prove the statement for continuous -theory. Consider a collection of dualizable categories. We need to show that We may assume that the categories are all equivalent: indeed, each is a retract in of the category In other words, we need to show that for a set and a dualizable category the map is an isomorphism. But this is the special case of 1) for the standard -structure on Indeed, by Proposition 1.105 the functor is a localizing invariant, and so is the functor When is compactly generated, the map is an isomorphism by Corollary 4.27.∎
@@
Put The functors are compatible with the functors hence we obtain a functor By construction, we have We need to show that in For each the inclusions define a fully faithful functor Its left adjoint is simply the structural functor Denote by the composition functors, Since the compositions vanish for we see that the direct sum of the functors is well-defined. Moreover, we have Now, denote by the composition, and put (where ). Equivalently, is the semi-orthogonal projection onto the left orthogonal to (where ). Denote by the natural transformations, Then Since the composition vanishes for we see that the direct sum is also a well-defined endofunctor of By the above, we obtain an exact triangle
@@
Since we may assume that Applying Proposition 4.21 to we obtain a short exact sequence
in Moreover, the source and the target of are compactly generated categories. Denote by the induced functor on the subcategories of compact objects. It suffices to show that is a -equivalence. As in the proof of proposition 4.21, we put and Note that the functor has an obvious section which corresponds to the collection of “representable” functors where is defined in (4.5). We need to show that in Choose any bijection Denote by the full stable subcategory generated by the objects Then is equivalent to the category where as usual denotes the poset Clearly, each inclusion has both left and right adjoints. Moreover, we see that the elements form a basis of It follows from Corollary 4.30 that Since we deduce that
@@
is also an isomorphism. By induction, we see that is an isomorphism for Denote by the induced map between the localizing invariants of dualizable categories. Arguing as above, we see that is an isomorphism for It remains to show that if is an isomorphism for some then is also an isomorphism. Let be a dualizable category. Then the functor is a strongly continuous localization (the right adjoint sends an object of to the corresponding constant sheaf). We denote by the kernel of Applying Proposition 4.31, we see that the composition
@@ 5.1. Continuous posets
Recall that a partially ordered set is called continuous if has directed supremums (joins), and the functor has a left adjoint Equivalently, this means the following. Assuming that has directed supremums, one defines another relation on namely, if for any directed system of elements of such that we have for some Then is continuous if and only if for each the poset is directed and we have equality
@@ 5.2. Sheaves of spectra on continuous posets
Let be a continuous poset. We denote by the category of sheaves on (where is considered as a category) for the above topology. This is in fact the category from Subsection 1.6. Namely, since every object of has the smallest covering sieve, a presheaf is a sheaf if and only if
@@ 5.3. Localizing invariants of categories of sheaves on continuous posets
Consider now a presheaf on with values in for We denote by the restriction functors for and by their right adjoints. Consider the category of -valued presheaves on (i.e. the category of sections of the associated cocartesian fibration over which is also cartesian). For we denote by the functor given by
The right adjoint to is given by Denote by the category of -valued sheaves on For we have a functor The right adjoint to is given by Our goal in this section is to prove the following result.
@@
1) The category is dualizable. 2) For each compact element the functor is strongly continuous and fully faithful. 3) The functors induce an isomorphism
@@
In particular, we have for This shows that for the functor commutes with coproducts. This proves 2). To prove 1), we first observe that the category is dualizable. Indeed, we have a -indexed semi-othogonal decomposition of the category in
Since are dualizable, so is the category by Proposition 1.81. Now, it follows from (5.2) that the sheafification functor has a left adjoint We deduce that the category is dualizable. To prove 3), consider the full subcategory By the above we have a short exact sequence in
where the functor is the left adjoint to the inclusion. We first observe the following. Claim 1. Let be a presheaf such that for any the map is zero. Then is contained in
@@
Next, we observe that the category has the following semi-orthogonal decomposition. Claim 2. For the functor is zero (resp. fully faithful) for (resp. for ). Moreover, we have a semi-orthogonal decomposition
@@
Indeed, it follows from (5.2) (by adjunction) that for the object is left orthogonal to On the other hand, it follows from Claim 1 that the object is contained in (since the latter presheaf can take non-zero values only for such that but ). This proves the isomorphism (5.5). We deduce that
@@
Hence, the adjunction unit is an isomorphism. We conclude that the functor is fully faithful for Similarly, if and then for This implies the semi-orthogonality. Finally, taking into account the vanishing of for we deduce that the categories generate This proves the claim.∎